### v1

# A Criterion for Perfectoid Purity and the Rationality of Thresholds

Abstract.
We introduce a new criterion providing a sufficient condition for a hypersurface in an unramified regular local ring to be perfectoid pure. The criterion is formulated in terms of an explicitly computable sequence of integers, called the splitting-order sequence. Our main theorem shows that if all entries of the sequence are at most p1p-1, then the hypersurface is perfectoid pure, and the perfectoid pure threshold can be computed explicitly from it. As a consequence, we prove that for any regular local ring $R$, the perfectoid pure threshold $\operatorname{ppt}(R,p)$ with respect to $p$ is always a rational number. Moreover, we show that for sufficiently large primes $p$, the cone over a Fermat type Calabi–Yau hypersurface is perfectoid pure, revealing new and unexpected examples of perfectoid pure singularities.
1. Introduction
The theory of singularities in mixed characteristic has recently undergone remarkable development, as seen in works such as [MSTWW], [BMPSTWW24], and [HLS]. This line of research has produced profound applications in both algebraic geometry and commutative algebra, including the mixed characteristic minimal model program [BMPSTWW23], [TY23], and Skoda-type theorems [HLS].
In this paper, we focus on the class of singularities known as perfectoid purity [p-pure], which can be regarded as an analogue of $F$-purity in positive characteristic. It is known that if $(R,\mathfrak{m})$ is a Noetherian local ring with $p\in\mathfrak{m}$, and if $R$ is a complete intersection such that $R/pR$ is quasi-$F$-split, then $R$ is perfectoid pure ([Yoshikawa25]). On the other hand, in positive characteristic, the celebrated Fedder-type criterion [kty] provides a practical method for testing quasi-$F$-purity for complete intersections, leading to a wealth of explicit examples.
In this paper, we establish a new numerical criterion that provides a sufficient condition for a hypersurface defined in an unramified regular local ring to be perfectoid pure. This criterion is expressed in terms of an explicitly computable sequence of integers, called the splitting–order sequence. This invariant generalizes the quasi-$F$-split height introduced in [Yoshikawa25], and provides both a concrete criterion for perfectoid purity and a $p$-adic expansion of the perfectoid pure threshold. As a corollary, we show that for any regular local ring $R$, the perfectoid pure threshold $\operatorname{ppt}(R,p)$ is always a rational number. This rationality result may be viewed as a partial analogue of [BMS]*Theorem 3.1, which proves that for a regular local ring of positive characteristic and any element $f$, the $F$-pure threshold $\operatorname{fpt}(R,f)$ is rational.
We now fix notation and recall the notion of perfectoid purity. Let $k$ be a perfect field and set
$A:=W(k)[[x_{1},\ldots,x_{N}]]$
with a Frobenius lift $\phi$ given by $\phi(x_{i})=x_{i}^{p}$ for $1\leq i\leq N$. For $a\in A$, set
$\Delta(a):=\frac{a^{p}-\phi(a)}{p}.$
Let $f\in A$ be such that $(p,f)$ forms a regular sequence. We say that $A/f$ is perfectoid pure if there exists a pure extension $A/f\to B$ to a perfectoid ring. If $A/f$ is perfectoid pure, the perfectoid pure threshold $\operatorname{ppt}(A/f,p)$ with respect to $p$ is defined as
[…]
see Proposition 2.2 for details. It is known [Yoshikawa25]*Theorem A that if $\overline{A/f}$ is quasi-$F$-split, then $A/f$ is perfectoid pure and $\operatorname{ppt}(A/f,p)>\frac{p-2}{p-1}$.
Next, we define the splitting–order sequence. Fix a generator $u\in\operatorname{Hom}_{\overline{A}}(F_{*}\overline{A},\overline{A})$ and set $\mathfrak{m}^{[p^{e}]}:=(p,x^{p^{e}}\mid x\in\mathfrak{m})$. For integers $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$, we define ideals $I(l_{1},\ldots,l_{n})$ of $\overline{A}$ inductively by
  • $I(l_{n}):=f^{p-l_{n}}\overline{A}$;
  • once $I(l_{2},\ldots,l_{n})$ is defined, set
    $I(l_{1},\ldots,l_{n}):=f^{p-l_{1}-1}\,u\bigl(F_{*}(\Delta(f)^{l_{1}}I(l_{2},\ldots,l_{n}))\bigr)+f^{p-l_{1}}\overline{A}.$
We then define the splitting–order sequence $\boldsymbol{s}(f)=(s_{0},s_{1},\ldots)$ of $f$ with values in $\{0,\ldots,p\}$ inductively as follows:
  • set $s_{0}:=0$;
  • once $s_{i}=p$ for some $i$, set $s_{n}=p$ for all $n>i$;
  • suppose $s_{0},\ldots,s_{n-1}$ have been defined with $s_{1},\ldots,s_{n-1}\leq p-1$, and define
    $s_{n}:=\max\bigl\{\,0\leq s\leq p\,\bigm|\,I(s_{1},\ldots,s_{n-1},s)\subseteq\mathfrak{m}^{[p]}\,\bigr\}.$
For alternative formulations of the splitting–order sequence, see Definition 3.6, and for the equivalence among them, see Theorem 5.2.
The splitting–order sequence serves as a perfectoid analogue of Fedder-type criteria in the theory of $F$-singularities. Our first main theorem shows that if all $s_{n}\leq p-1$, then $A/f$ is perfectoid pure, and moreover it expresses $\operatorname{ppt}(A/f,p)$ explicitly in terms of $\boldsymbol{s}(f)$.
Theorem A.
(cf. Theorem 4.5) In the above setting, if $s_{n}\leq p-1$ for every $n\geq 1$, then $A/f$ is perfectoid pure with
$\operatorname{ppt}(A/f,p)=\sum_{n\geq 1}\frac{p-1-s_{n}}{p^{n}}.$
An important open problem is whether the converse direction also holds.
Question 1.1.
Does the converse direction of Theorem A hold? That is, if $A/f$ is perfectoid pure, must we have $s_{n}\leq p-1$ for all $n\geq 1$?
We obtain a partial answer to this question in the following theorem, which completely resolves it when $p=2$.
Theorem B.
(cf. Theorem 4.7) In the above setting, we have:
  1. (1)
    If $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)\geq\frac{p-2}{p-1}$, then $s_{n}\leq 1$ for every $n\geq 1$.
  2. (2)
    If $s_{1}=\cdots=s_{r}=p-1$ and $s_{r+1}=p$ for some $r\geq 2$, then $A/f$ is not perfectoid pure.
  3. (3)
    If $A/f$ is regular, then $s_{n}\leq p-1$ for every $n\geq 0$.
Combining Theorems A and B, we obtain the following fundamental result, which shows that $\operatorname{ppt}(R,p)$ is always a rational number for a regular local ring $R$.
Theorem C.
(Corollary 5.4) Let $(R,\mathfrak{m})$ be a regular local ring with $0\neq p\in\mathfrak{m}$ such that $R/pR$ is $F$-finite. Then $\operatorname{ppt}(R,p)$ is a rational number.
Finally, in Section 6 we compute explicit examples that illustrate these results. In particular, we show that perfectoid purity appears in a surprisingly broad range of situations, including cones over Fermat-type Calabi–Yau hypersurfaces.
Example 1.2 (Theorems 6.2, 6.7 and 6.6).
  1. (1)
    Let $f=x_{1}^{N}+\cdots+x_{p}^{N}$. If $p>N$, then $A/f$ is perfectoid pure.
  2. (2)
    Let $N=4$, $p\equiv 3\pmod{4}$, and $f=x_{1}^{4}+\cdots+x_{4}^{4}$. Then $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=2/(p^{2}-1)$. In particular, the lift of the Fermat-type K3 cone for $p=3$ is perfectoid pure; such a K3 surface has Artin invariant $1$.
  3. (3)
    Let $p=2$, $N=4$, and
    $f=x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+x_{4}^{4}+x_{1}^{2}x_{2}^{2}+x_{1}^{2}x_{3}^{2}+x_{2}^{2}x_{3}^{2}+x_{1}x_{2}x_{3}(x_{1}+x_{2}+x_{3})+px_{1}x_{2}x_{3}x_{4},$
    then $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$. Moreover, $\mathrm{Proj}(\overline{A/f})$ is an RDP K3 surface of Artin invariant $1$ by [DK03]*Theorem 1.1(vii).
These computations show that perfectoid purity can occur even in unexpected geometric settings, highlighting the richness of this class of singularities. Our approach therefore provides a new and effective computational framework for studying perfectoid analogues of $F$-singularity invariants.
Acknowledgements.
The author wishes to express his gratitude to Teppei Takamastu and Linquan Ma for valuable discussion. The author was supported by JSPS KAKENHI Grant number JP24K16889.
2. Preliminaries
In this section, we collect the basic definitions and notations used throughout the paper. We recall the conventions for $\delta$-rings, Frobenius lifts, and perfectoid structures, which will serve as the foundation for later sections.
2.1. Notation and conventions
In this subsection, we summarize the notation and terminology used throughout this paper.
  1. (1)
    We fix a prime number $p$ and set $\mathbb{F}_{p}:=\mathbb{Z}/p\mathbb{Z}$.
  2. (2)
    For a ring $R$, we set $\overline{R}:=R/pR$.
  3. (3)
    For a ring $R$ of characteristic $p>0$, we denote by $F\colon R\to R$ the absolute Frobenius ring homomorphism, defined by $F(r)=r^{p}$ for all $r\in R$.
  4. (4)
    For a ring $R$ of characteristic $p>0$, we say that $R$ is $F$-finite if $F$ is finite.
  5. (5)
    For a ring $R$ of characteristic $p>0$, we define the perfection of $R$ as $R_{\mathrm{perf}}:=\mathrm{colim}_{F}R$.
  6. (6)
    We use terminology $\delta$-ring as defined in [BS]*Definition 2.1. For a $\delta$-ring $A$, the delta map is denoted by $\delta$ and set $\Delta:=-\delta$.
  7. (7)
    We use the terminology perfectoid as defined in [BMS19]*Definition 3.5. For a semiperfectoid ring $R$, its perfectoidization is denoted by $R_{\mathrm{perfd}}$ (cf. [BS]*Section 7).
  8. (8)
    Let $R$ be a ring and $f$ an element of $R$. A sequence $\{f^{1/p^{e}}\}_{e\geq 0}$ is called a compatible system of $p$-power roots of $f$ if $(f^{1/p^{e+1}})^{p}=f^{1/p^{e}}$ for all $e\geq 0$ and $f^{1/p^{0}}=f$. Furthermore, the ideal $(f^{1/p^{e}}\mid e\geq 0)$ of $R$ is denoted by $(f^{1/p^{\infty}})$.
  9. (9)
    For a ring $R$ and a ring homomorphism $\phi\colon R\to R$, we say that $\phi$ is a Frobenius lift if $\phi(a)\equiv a^{p}\ \textrm{mod}\ p$ for every $a\in R$.
  10. (10)
    For a ring homomorphism $f\colon A\to B$, we say that $f$ is pure if for every $A$-module $M$, the homomorphism $M\to M\otimes_{A}B$ is injective.
2.2. Frobenius lifts and perfectoid covers
We construct certain perfectoid covers associated with a Frobenius lift on a regular local ring. This technical result will later be used to connect the splitting-order sequence to perfectoid purity.
Proposition 2.1.
Let $(A,\mathfrak{m})$ be a complete regular local ring with $0\neq p\in\mathfrak{m}$. Then $A$ is unramified if and only if $A$ admits a Frobenius lift.
Proof.
First, assume that $A$ admits a Frobenius lift $\phi$. Since $A$ is regular, $\phi$ is flat. Reducing modulo $p$, we obtain that the Frobenius map
$F\colon\overline{A}\longrightarrow F_{*}\overline{A}$
is flat. By Kunz’s theorem, this implies that $\overline{A}$ is regular. Hence $A$ is unramified.
Conversely, assume that $A$ is unramified. Then, by [Matsumura]*Theorem 29.7, we have
$A\simeq C[[x_{1},\ldots,x_{N}]]$
for some complete discrete valuation ring $C$ with uniformizer $p$. By [Matsumura]*Theorem 29.2, the ring $C$ admits a Frobenius lift, and therefore so does $A$. ∎
Proposition 2.2.
Let $(A,\mathfrak{m})$ be a regular local ring with a finite Frobenius lift $\phi$, and assume $0\neq p\in\mathfrak{m}$. Let $f\in A$. Then there exist a morphism of $\delta$-rings $A\to B$ to a perfect prism $(B,(d))$ such that:
  • $A\to B/(d)$ is faithfully flat;
  • $B/(d)$ contains compatible systems of $p$-power roots of $f$ and $p$;
  • we have
    $\operatorname{ppt}(A/f,p)=\sup\bigl\{\alpha\in\mathbb{Z}[1/p]_{\geq 0}\,\big|\,A/f\to B/(d,f^{1/p^{\infty}})\xrightarrow{\cdot p^{\alpha}}B/(d,f^{1/p^{\infty}})\text{ is pure}\bigr\}.$
Proof.
Set $A^{\prime}:=A[X_{0}]$ and define a ring homomorphism $\phi^{\prime}\colon A^{\prime}\to A^{\prime}$ by $\phi^{\prime}|_{A}=\phi$ and $\phi^{\prime}(X_{0})=X_{0}^{p}$; then $A\to A^{\prime}$ is a morphism of $\delta$-rings. Define
[…]
so that $(B^{\prime},(d:=X_{0}-p))$ is a perfect prism. By construction, $A\to B^{\prime}$ is a morphism of $\delta$-rings, and $A\to B^{\prime}/(d)$ is $p$-completely faithfully flat.
By André’s flatness lemma, there exists a $(p,d)$-faithfully flat morphism of perfect prisms $(B^{\prime},(d))\to(B,(d))$ such that $B/(d)$ contains compatible systems of $p$-power roots of $f$ and $p$. Hence $A\to B/(d)$ is also $p$-completely faithfully flat. By [Bhatt20]*Lemma 5.15, this implies that $A\to B/(d)$ is faithfully flat.
Finally, we verify the third condition. Let $C$ be a perfectoid $A/f$-algebra. By the proof of [Yoshikawa25]*Proposition 2.9, it suffices to show that, after replacing $C$ by a $p$-completely faithfully flat extension, there exists an $A$-algebra homomorphism $B/(d)\to C$. After replacing $C$ by an $p$-completely faithfully flat extension, we may assume that $C$ contains a root of every monic polynomial. We define an $A$-algebra homomorphism $A^{\prime}\to C$ by $X_{0}\mapsto p$. Since $\phi^{\prime}$ is finite, there exists an $A$-algebra homomorphism $B^{\prime}\to C$ such that $d\mapsto 0$. Then the natural map $C\to C^{\prime}:=C\widehat{\otimes}_{B^{\prime}/(d)}B/(d)$ is $p$-completely faithfully flat, and there is an induced $A$-algebra homomorphism $B\to C^{\prime}$, as desired. ∎
3. Splitting-order sequences
In this section, we introduce the notion of a splitting-order sequence, a numerical invariant encoding how far an element is from being Frobenius split. We establish its basic properties and functoriality, and relate it to quasi-$F$-splitting.
Notation 3.1.
Let $A$ be a $\delta$-ring, and set $\Delta:=-\delta$. For $f\in A$ and integers $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$, we define an $\overline{A}$-module $Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ by
$Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}:=\frac{\overline{A/f^{\,p}}\oplus\cdots\oplus F^{n-1}_{*}\overline{A/f^{p}}}{\Bigl((a_{1}f^{\,l_{1}},-a_{1}^{p}\Delta(f)^{\,l_{1}}+a_{2}f^{l_{2}},\ldots,-a_{n-1}\Delta(f)^{l_{n-1}}+a_{n}f^{l_{n}})\ \Bigm|\ a_{1},\ldots,a_{n}\in\overline{A}\Bigr)}$
We also define $\overline{A}$-module homomorphisms
$\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}\colon\overline{A/f^{\,l_{1}+1}}\longrightarrow Q_{A,f,\left(l_{1},\ldots,l_{n}\right)},\qquad V\colon F^{\,n-1}_{*}\overline{A/f^{\,l_{n}}}\longrightarrow Q_{A,f,\left(l_{1},\ldots,l_{n}\right)},$
as follows.
  • Consider the $\overline{A}$-linear map $\overline{A}\to Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$, $a\mapsto(a,0,\ldots,0)$. In $Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ we have
    $(f^{\,l_{1}+1},0,\ldots,0)=(0,f^{\,p}\Delta(f)^{\,l_{1}},0,\ldots,0)=0$
    by the defining relations. Hence the map factors through $\overline{A/f^{\,l_{1}+1}}$, yielding
    $\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}\colon\overline{A/f^{\,l_{1}+1}}\longrightarrow Q_{A,f,\left(l_{1},\ldots,l_{n}\right)},\qquad a\longmapsto(a,0,\ldots,0).$
  • Similarly, the $\overline{A}$-linear map $F^{\,n-1}_{*}\overline{A}\to Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$, $a\mapsto(0,\ldots,0,a)$, factors through $F^{\,n-1}_{*}\overline{A/f^{\,l_{n}}}$, giving
    $V\colon F^{\,n-1}_{*}\overline{A/f^{\,l_{n}}}\longrightarrow Q_{A,f,\left(l_{1},\ldots,l_{n}\right)},\qquad a\longmapsto(0,\ldots,0,a).$
When the base ring $A$ is clear from the context, we simply write $Q_{f,\left(l_{1},\ldots,l_{n}\right)}$ and $\Phi_{f,\left(l_{1},\ldots,l_{n}\right)}$ in place of $Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}$ and $\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}$, respectively.
Proposition 3.2.
With notation as in 3.1, fix $f,f^{\prime}\in A$, and integers $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$. If $f\equiv f^{\prime}\pmod{p^{2}}$, then
$Q_{f,\left(l_{1},\ldots,l_{n}\right)}\;\simeq\;Q_{f^{\prime},\left(l_{1},\ldots,l_{n}\right)}.$
Proof.
It suffices to show that $\Delta(f)\equiv\Delta(f^{\prime})\pmod{p}$. Write $f-f^{\prime}=p^{2}g$ for some $g\in A$. Then
$\Delta(f)=\Delta(f^{\prime}+p^{2}g)\equiv\Delta(f^{\prime})+\Delta(p^{2}g)\equiv\Delta(f^{\prime})\pmod{p}$,
as desired. ∎
Proposition 3.3.
With notation as in 3.1, fix $f\in A$. Then there exists an $\overline{A}$-module isomorphism
$\overline{W}_{n}(A/f)\;\xrightarrow{\ \simeq\ }\;Q_{f,\left(1,\ldots,1\right)},$
where the $\overline{A}$-module structure on the left-hand side is given via the ring homomorphism $s_{\phi}\colon A\to W_{n}(A)$ (cf. [Yoshikawa25]*Proposition 3.2).
Proof.
By [Yoshikawa25]*Theorem 3.7, there is an $A$-module isomorphism
$\Psi\colon\overline{W}_{n}(A)\xrightarrow{\ \simeq\ }\overline{A}\oplus F_{*}\overline{A}\oplus\cdots\oplus F^{n-1}_{*}\overline{A},$
defined by
$\Psi(a_{0},\ldots,a_{n-1})=\bigl(a_{0},\,\Delta_{1}(a_{0})+a_{1},\,\ldots,\,\Delta_{n-1}(a_{0})+\cdots+a_{n-1}\bigr).$
We define a map
$\Psi^{\prime}\colon\overline{W}_{n}(A)\longrightarrow Q_{f,\left(1,\ldots,1\right)}$
by
$\Psi^{\prime}(a_{0},\ldots,a_{n-1})=\bigl(a_{0},\,-(\Delta_{1}(a_{0})+a_{1}),\,\ldots,\,(-1)^{n-1}(\Delta_{n-1}(a_{0})+\cdots+a_{n-1})\bigr).$
We show that $\Psi^{\prime}$ induces an isomorphism after passing to the quotient by $W_{n}(f)$.
For any $(a_{0},\ldots,a_{n-1})\in\overline{W}_{n}(A)$, we compute:
$\Psi^{\prime}(a_{0}f,a_{1}f,\ldots,a_{n-1}f)$ $=\bigl(a_{0}f,\,-a_{0}^{p}\Delta_{1}(f)-a_{1}f,\,\ldots,\,(-1)^{n-1}(a_{0}^{p^{n-1}}\Delta_{n-1}(f)+\cdots+a_{n-1}f)\bigr)$
$\overset{(\star_{1})}{=}\bigl(a_{0}f,\,-a_{0}^{p}\Delta_{1}(f)-a_{1}f,\,\ldots,\,(-1)^{n-1}(a_{n-1}^{p}\Delta_{1}(f)+a_{n}f)\bigr)$
$\overset{(\star_{2})}{=}(0,\ldots,0),$
where $(\star_{1})$ follows from [Yoshikawa25]*Theorem 3.6 and $(\star_{2})$ from the definition of $Q_{f,\left(1,\ldots,1\right)}$.
Hence $\Psi^{\prime\prime}$ induces an $\overline{A}$-module homomorphism
$\Psi^{\prime\prime}\colon\overline{W}_{n}(A/f)\longrightarrow Q_{f,\left(1,\ldots,1\right)}.$
By construction, $\Psi^{\prime}$ is surjective, so $\Psi^{\prime\prime}$ is surjective as well.
To prove injectivity, let $\alpha=(a_{0},\ldots,a_{n-1})\in\overline{W}_{n}(A)$ satisfy $\Psi^{\prime}(\alpha)=0$. Then $a_{0}\in(f)$. Replacing $a_{0}$ by zero, we proceed inductively to see that each $a_{i}\in(f)$ for all $i$. Therefore $\alpha\in\overline{W}_{n}(fA)$, and hence $\Psi^{\prime\prime}$ is injective.
Thus $\Psi^{\prime\prime}$ is an isomorphism of $\overline{A}$-modules, as claimed. ∎
Proposition 3.4.
With notation as in 3.1, let $\varphi\colon A\to B$ be a morphism of $\delta$-rings. Take $f\in A$ and $g\in B$ such that $\varphi(f)\in(g)$. Then, for all integers $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$, the map $\varphi$ induces an $\overline{A}$-module homomorphism
$Q_{f,\left(l_{1},\ldots,l_{n}\right)}\;\longrightarrow\;Q_{g,\left(l_{1},\ldots,l_{n}\right)}.$
Proof.
It suffices to show that, for $0\leq l\leq p-1$, one has
$(\varphi(f^{\,l}),-\varphi(\Delta(f)^{\,l}))\in\bigl((bg^{\,l},-b^{p}\Delta(g)^{\,l})\ \bigm|\ b\in\overline{A}\bigr)\subseteq\overline{B/g^{p}}\oplus F_{*}\overline{B/g^{p}}.$
Write $\varphi(f)=gh$. Then, since $\varphi$ is a morphism of $\delta$-rings,
$(\varphi(f^{\,l}),-\varphi(\Delta(f)^{\,l}))=(h^{l}g^{l},-\Delta(gh)^{l})=(h^{l}g^{l},-h^{pl}\Delta(g)^{l}),$
where the last equality holds because $(0,-g^{p}\Delta(h))=0$ in $\overline{B/g^{p}}\oplus F_{*}\overline{B/g^{p}}$. ∎
Proposition 3.5.
With notation as in 3.1, fix $f\in A$. Assume that $p,f$ form a regular sequence. For $n\geq 2$, $0\leq l_{1},\ldots,l_{n-1}\leq p-1$, and $0\leq l_{n}\leq p$, there is an exact sequence
[…]
where the map $R$ is given by
$R(a_{1},\ldots,a_{n})=(a_{1},\ldots,a_{n-1}).$
Proof.
It suffices to show that $V$ is injective. Take $a\in\overline{A}$ such that $(0,\ldots,0,a)=0$ in $Q_{f,\left(l_{1},\ldots,l_{n}\right)}$. Then there exist $a_{1},\ldots,a_{n}\in\overline{A}$ satisfying
$(0,\ldots,0,a)=(a_{1}f^{l_{1}},\,-a_{1}^{p}\Delta(f)^{l_{1}}+a_{2}f^{l_{2}},\,\ldots,\,-a_{n-1}^{p}\Delta(f)^{l_{n-1}}+a_{n}f^{l_{n}})$
in $\overline{A/f^{p}}\oplus\cdots\oplus F^{n-1}_{*}\overline{A/f^{p}}$.
Since $p,f$ form a regular sequence, we have $a_{1}\in f^{p-l_{1}}\overline{A}$. Because $l_{1}\leq p-1$, it follows that $a_{1}^{p}\in f^{p}\overline{A}$; in particular, $a_{2}f^{l_{2}}=0$ in $\overline{A/f^{p}}$ when $n\geq 3$. Repeating this argument inductively, we obtain $a=a_{n}f^{l_{n}}$ in $\overline{A/f^{p}}$, hence $a\in f^{l_{n}}\overline{A}$, as desired. ∎
Definition 3.6.
With notation as in 3.1, fix $f\in A$. We define a splitting-order sequence $\boldsymbol{s}(f)=(s_{i})_{i\geq 0}$ of $f$ with values in $\{0,\ldots,p\}$ inductively as follows:
  • Set $s_{0}:=0$.
  • Once $s_{i}=p$ for some $i$, set $s_{n}=p$ for all $n>i$.
  • Suppose $s_{0},\ldots,s_{n-1}$ have been defined for some $n\geq 1$ with $s_{0},\ldots,s_{n-1}\leq p-1$. We define
[…]
Remark 3.7.
  • If $\Phi_{f,\left(s_{0},\ldots,s_{n-1}\right)}$ is not pure, then $\Phi_{f,\left(s_{0},\ldots,s_{n-1},0\right)}$ is also not pure since $Q_{f,\left(s_{0},\ldots,s_{n-1},0\right)}\simeq Q_{f,\left(s_{0},\ldots,s_{n-1}\right)}$. In particular, the set in Definition 3.6 is non-empty.
  • In the setting of Definition 3.6, the splitting-order sequence is invariant modulo $p^{2}$; indeed, if $f\equiv f^{\prime}\pmod{p^{2}}$, then $\boldsymbol{s}(f)=\boldsymbol{s}(f^{\prime})$ by Proposition 3.2.
Proposition 3.8.
Let $(A,\mathfrak{m})$ be a Noetherian local ring with a finite Frobenius lift $\phi$, and assume that $p\in\mathfrak{m}$ and $A$ is $p$-torsion free. Let $f\in A$ and let $\boldsymbol{s}(f)$ denote the splitting–order sequence of $f$. Then $A/f$ is quasi-$F$-split of height $h$ if and only if $s_{1}=\cdots=s_{h-1}=1$ and $s_{h}=0$.
Proof.
We first observe that there is an isomorphism
$Q_{f,\left(s_{0},\ldots,s_{n}\right)}\xrightarrow{\ \simeq\ }F_{*}Q_{f,\left(s_{1},\ldots,s_{n}\right)},\qquad(a_{0},\ldots,a_{n})\longmapsto(a_{0}^{p}+a_{1},\ldots,a_{n}).$
Composing $\Phi_{f,\left(s_{0},\ldots,s_{n}\right)}$ with the above map yields
$\overline{A/f}\longrightarrow F_{*}Q_{f,\left(s_{1},\ldots,s_{n}\right)},\qquad a\longmapsto(a^{p},0,\ldots,0),$
which agrees with the morphism $\Phi_{A/f,n}$ after composing with the isomorphism in Proposition 3.3. Therefore, the claim follows directly from the definitions of the sequence $\boldsymbol{s}(f)$ and quasi-$F$-splitting. ∎
Proposition 3.9.
Let $(A,\mathfrak{m})$ be a Noetherian local ring with a finite Frobenius lift $\phi$, and assume that $p\in\mathfrak{m}$. Let $\iota\colon A\to\widehat{A}$ denote the $\mathfrak{m}$-adic completion. Then $\phi$ induces a Frobenius lift $\widehat{\phi}$ on $\widehat{A}$. We equip $A$ and $\widehat{A}$ with the $\delta$-ring structures corresponding to $\phi$ and $\widehat{\phi}$, respectively. Take $f\in A$ such that $p,f$ form a regular sequence. Then $\boldsymbol{s}(f)=\boldsymbol{s}(\iota(f))$.
Proof.
Since $\overline{A}$ is $F$-finite, we have the following commutative diagram:
${\bigl(\overline{A/f}\otimes_{\overline{A}}\overline{\widehat{A}}\bigr)}$${\bigl(Q_{A,f,\left(l_{1},\ldots,l_{n}\right)}\otimes_{\overline{A}}\overline{\widehat{A}}\bigr)}$${\overline{\widehat{A}/\iota(f)}}$${Q_{\widehat{A},\iota(f),\left(l_{1},\ldots,l_{n}\right)}}$$\scriptstyle{\Phi_{A,f,\left(l_{1},\ldots,l_{n}\right)}\otimes\overline{\widehat{A}}}\scriptstyle{\simeq}\scriptstyle{\simeq}$$\scriptstyle{\Phi_{\widehat{A},\iota(f),\left(l_{1},\ldots,l_{n}\right)}}$
for $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$. The desired assertion follows immediately from the diagram. ∎
4. Perfectoid purity versus splitting-order sequences
This section establishes the relationship between perfectoid purity and the boundedness of the splitting-order sequence. The key result shows that the perfectoid $p$-purity threshold can be expressed in terms of the splitting-order sequence.
Lemma 4.1.
With notation as in 3.1, fix $f\in A$, $n\geq 1$, $0\leq l_{1},\ldots,l_{n-1}\leq p-1$, and $0\leq l_{n}\leq p$. Assume that $(A,\mathfrak{m})$ is a Noetherian local ring such that $p,f$ form a regular sequence. Suppose further that $A$ is Cohen–Macaulay. Then $H^{i}_{\mathfrak{m}}(Q_{f,\left(l_{1},\ldots,l_{n}\right)})=0$ for all $i<d:=\dim(\overline{A/f})$, and the morphism
$H^{d}_{\mathfrak{m}}(F^{n-1}_{*}\overline{A/f^{l_{n}}})\longrightarrow H^{d}_{\mathfrak{m}}(Q_{f,\left(l_{1},\ldots,l_{n}\right)})$
induced by $V$ is injective.
Proof.
Since $A$ is Cohen–Macaulay and $p,f$ form a regular sequence, each $\overline{A/f^{l}}$ is Cohen–Macaulay for $l\geq 1$. The desired vanishing and injectivity then follow from the exact sequence in Proposition 3.5. ∎
Proposition 4.2.
Let $(A,(d))$ be a perfect prism, and set $T:=\overline{d}\in\overline{A}$, then we obtain the isomorphism
$\overline{A}/(T^{l_{0}+l_{1}/p+\cdots+l_{n}/p^{n}})\simeq Q_{d,\left(l_{0},l_{1},\ldots,l_{n}\right)}\ \,a\mapsto(a,0,\cdots,0)$
for $0\leq l_{0},\ldots,l_{n-1}$ and $0\leq l_{n}\leq p$.
Proof.
We prove the assertion by induction on $n$. If $n=0$, it is clear. For $n\geq 1$, we have
$Q_{d,\left(l_{0},\ldots,l_{n}\right)}\simeq\overline{A}/T^{l_{0}}\oplus F_{*}\overline{A}/(T^{l_{1}+\cdots+l_{n}/p^{n-1}})/\{(aT^{l_{0}},-a^{p}\Delta(d)^{l_{0}})\mid a\in\overline{A}\}=:P.$
First, the homomorphism
$\overline{A}/(T^{l_{0}+l_{1}/p+\cdots+l_{n}/p^{n}})\to P$
is well-defined, indeed, we have
$(T^{l_{0}+l_{1}/p+\cdots+l_{n}/p^{n}},0)=(0,T^{l_{1}+\cdots+l_{n}/p^{n-1}}\Delta(d)^{l_{0}})=0.$
Furthermore, the inverse homomorphism
$P\to\overline{A}/(T^{l_{0}+l_{1}/p+\cdots+l_{n}/p^{n}})$
is defined by
$(a,b)\mapsto a-F^{-1}(b)\Delta(d)^{-l_{0}}T^{l_{0}},$
then it is well-defined, as desired. ∎
Lemma 4.3.
With notation as in 3.1, fix $f\in A$, $n\geq 1$, $0\leq l_{1},\ldots,l_{n-1}\leq p-1$, and set $l_{n}:=p$. Let $(B,(d))$ be a perfect prism with Frobenius lift $\phi$, and let $\varphi\colon A\to B$ be a morphism of $\delta$-rings. Assume there exists $y\in B$ such that $\phi(y)=y^{p}$ and $y-f\in d\overline{B}$. Then for every element
$\alpha\in\operatorname{Ker}\!\Bigl(\overline{A}\oplus F_{*}\overline{A}\oplus\cdots\oplus F^{n}_{*}\overline{A}\longrightarrow Q_{A,f,\left(0,l_{1},\ldots,l_{n}\right)}\Bigr),$
we have
$\alpha\,y^{1-1/p^{n-1}}\in f\,Q_{B,d,\left(0,l_{1},\ldots,l_{n-1},1\right)}.$
Proof.
There exist elements $a_{0},\ldots,a_{n}\in\overline{A}$ such that
$\alpha=(a_{0},\,-a_{0}^{p}+a_{1}f^{l_{1}},\,-a_{1}^{p}\Delta(f)^{l_{1}}+a_{2}f^{l_{2}},\,\ldots,\,-a_{n-1}^{p}\Delta(f)^{l_{n-1}}+a_{n}f^{p}).$
We first note that
$(0,\ldots,0,a_{n}f^{p})\,y^{1-1/p^{n-1}}=(0,\ldots,0,a_{n}f^{p}y^{p^{n}-p})=(0,\ldots,0,a_{n}f^{p^{n}})$
in $Q_{B,d,\left(0,l_{1},\ldots,l_{n-1},1\right)}$. Hence we may assume $a_{n}=0$. If $n=1$, then
$\alpha=(a_{0},-a_{0}^{p})=0\text{ in }Q_{B,f,\left(0,1\right)},$
as desired. Thus we may further assume $n\geq 2$.
Next, we reduce to the case $a_{n-1}=0$. Consider the $\overline{B}$-module homomorphism
$F^{n-1}_{*}Q_{B,d,\left(l_{n-1},1\right)}\to Q_{B,d,\left(0,l_{1},\ldots,l_{n-1},1\right)},\qquad(a,b)\mapsto(0,\ldots,0,a,b).$
Let
$\beta:=(a_{n-1}f^{l_{n-1}},-a_{n-1}^{p}\Delta(f)^{l_{n-1}})y^{p^{n-1}-1}\in Q_{B,d,\left(l_{n-1},1\right)}.$
Write $f\equiv y+ud\pmod{pB}$ for some $u\in B$. Then
$\Delta(f)\equiv u^{p}\Delta(d)\pmod{d\overline{B}}$.
Thus
$\beta=(a_{n-1}f^{l_{n-1}},-(a_{n-1}u^{l_{n-1}})^{p}\Delta(d)^{l_{n-1}})y^{p^{n-1}-1}=(a_{n-1}y^{p^{n-1}-1}(f^{l_{n-1}}-(ud)^{l_{n-1}}),0)$
in $Q_{B,d,\left(l_{n-1},1\right)}$. Since $f^{l_{n-1}}-(ud)^{l_{n-1}}\in y\overline{B}$, we have $a_{n-1}y^{p^{n-1}-1}(f^{l_{n-1}}-(ud)^{l_{n-1}})\in y^{p^{n-1}}\overline{B}$. Because $n-1\geq 1$, this implies $\beta\in f^{p^{n-1}}Q_{B,d,\left(l_{n-1},1\right)}$, and hence
$(0,\ldots,0,a_{n-1}f^{l_{n-1}},-a_{n-1}^{p}\Delta(f)^{l_{n-1}})\in f\,Q_{B,d,\left(0,l_{1},\ldots,l_{n-1},1\right)}.$
Thus we may assume $a_{n-1}=0$.
Next, set $l_{0}=0$. For each $0\leq i\leq n-2$, define a $\overline{B}$-linear homomorphism
$\psi_{i}\colon F^{i}_{*}Q_{B,d,\left(l_{i},p\right)}\longrightarrow Q_{B,d,\left(0,l_{1},\ldots,l_{n}\right)}$
as follows: for $(a,b)\in F^{i}_{*}Q_{B,d,\left(l_{i},p\right)}$, let $\psi_{i}(a,b)=(c_{0},\ldots,c_{n})$, where
$c_{j}=\begin{cases}a&\text{if }j=i,\\ b&\text{if }j=i+1,\\ 0&\text{otherwise.}\end{cases}$
It suffices to show that
$\beta_{i}:=(a_{i}f^{l_{i}},-a_{i}^{p}\Delta(f)^{l_{i}})\,y^{p^{i}-1/p^{n-i}}\in f^{p^{i}}Q_{B,d,\left(l_{i},p\right)}\quad\text{for all }0\leq i\leq n-1.$
If $i=0$, then $\beta_{0}=0$ in $Q_{B,d,\left(l_{0},p\right)}$ since $l_{0}=0$.
Now assume $n-2\geq i\geq 1$. Then
$\Delta(f)\equiv u^{p}\Delta(d)\pmod{(y,d^{p})\overline{B}}$.
It follows that
$a_{i}^{p}y^{p^{i+1}-1/p^{n-i-2}}\Delta(f)^{l_{i}}\equiv(a_{i}y^{p^{i}-1/p^{n-i-1}}u^{l_{i}})^{p}\Delta(d)^{l_{i}}\pmod{(y^{p^{i+1}},d^{p})\overline{B}}$,
since $n-i-2\geq 0$. Moreover, as $i+1\geq 1$, we have $y^{p^{i+1}}\equiv f^{p^{i+1}}\pmod{d^{p}\overline{B}}$, and therefore
$\beta_{i}\equiv(a_{i}f^{l_{i}},-(a_{i}u^{l_{i}})^{p}\Delta(d)^{l_{i}})\,y^{p^{i}-1/p^{n-i-1}}\pmod{f^{p^{i}}Q_{B,d,\left(l_{i},p\right)}}$.
Hence
$\beta_{i}\equiv(a_{i}f^{l_{i}}-a_{i}(ud)^{l_{i}},0)\,y^{p^{i}-1/p^{n-i-1}}\pmod{f^{p^{i}}Q_{B,d,\left(l_{i},p\right)}}$.
Since $f^{l_{i}}-(ud)^{l_{i}}\in y\overline{B}$, we have $a_{i}y^{p^{i}-1/p^{n-i-1}}(f^{l_{i}}-(ud)^{l_{i}})\in y^{p^{i}}\overline{B}$. Because $i\geq 1$, this implies $\beta_{i}\in f^{p^{i}}Q_{B,d,\left(l_{i},p\right)}$, as desired. ∎
Lemma 4.4.
Let $(A,\mathfrak{m})$ be a regular local ring equipped with a finite Frobenius lift $\phi$, and assume that $0\neq p\in\mathfrak{m}$. We endow $A$ with the $\delta$-ring structure induced by $\phi$. Let $f\in A$. Then there exist a perfect prism $(B,(d))$, a morphism of $\delta$-rings $\varphi\colon A\to B$, and an element $y\in B$ such that
$\phi(y)=y^{p},\qquad y-f\in(d),\qquad\text{and}\qquad A\to B/(d)\text{ is faithfully flat.}$
In particular, $B/(d)$ is a Cohen–Macaulay $A$-algebra and $A\to B/(d)$ is pure.
Proof.
Take a perfect prism $(B,(d))$ as constructed in Proposition 2.2; then $A\to B/(d)$ is faithfully flat. Define $y\in B$ to be the element corresponding to $f^{\flat}]\in W((B/(d))^{\flat})$ under the canonical isomorphism $W((B/(d))^{\flat})\simeq B$. By construction we have $\phi(y)=y^{p}$ and $y-f\in(d)$, as desired. ∎
Theorem 4.5.
(Theorem A) Let $(A,\mathfrak{m})$ be a regular local ring with a finite Frobenius lift $\phi$ and $p\in\mathfrak{m}$. Fix the $\delta$-ring structure on $A$ induced by $\phi$. Let $f\in A$ be such that $p,f$ is a regular sequence. Let $\boldsymbol{s}(f)=(s_{i})_{i\geq 0}$ be the splitting-order sequence of $f$. Assume that $s_{i}\leq p-1$ for every $i\geq 0$. Then:
  1. (1)
    $A/f$ is perfectoid pure.
  2. (2)
    $\operatorname{ppt}(A/f,p)=\sum_{i\geq 1}\frac{p-1-s_{i}}{p^{i}}$.
Proof.
First, we prove (1). Set $d:=\dim(\overline{A/f})$ and take a nonzero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{A/f})$ in the socle. Fix $n\geq 1$. By the definition of $\boldsymbol{s}(f)$, the map $\Phi_{f,\left(s_{0},\ldots,s_{n-1}\right)}$ is not pure, while $\Phi_{f,\left(s_{0},\ldots,s_{n-1},s_{n}+1\right)}$ is pure; in particular $s_{n}+1\leq p$.
By Proposition 3.5 there exists $\eta^{\prime}_{n}\in H^{d}_{\mathfrak{m}}(\overline{A/f^{\,s_{n}+1}})$ with
$V(\eta^{\prime}_{n})=\Phi_{f,\left(s_{0},\ldots,s_{n-1},s_{n}+1\right)}(\eta).$
Since $\Phi_{f,\left(s_{0},\ldots,s_{n-1},s_{n}\right)}(\eta)=0$, Lemma 4.1 yields
$\bigl(H^{d}_{\mathfrak{m}}(\overline{A/f^{\,s_{n}+1}})\!\to H^{d}_{\mathfrak{m}}(\overline{A/f^{\,s_{n}}})\bigr)(\eta^{\prime}_{n})=0,$
so there exists $\eta_{n}\in H^{d}_{\mathfrak{m}}(\overline{A/f})$ with $f^{s_{n}}\eta_{n}=\eta^{\prime}_{n}$.
Write $\tau:=\eta/f\in H^{d+1}_{\mathfrak{m}}(\overline{A})$ and $\tau_{n}:=\eta_{n}/f\in H^{d+1}_{\mathfrak{m}}(\overline{A})$. Choose a regular sequence $x_{1},\ldots,x_{d}\in A$ so that $x_{1},\ldots,x_{d},f$ is a system of parameters, and elements $a,b\in\overline{A}$ with
$\tau=\Bigl[\frac{a}{x_{1}\cdots x_{d}f}\Bigr],\qquad\tau_{n}=\Bigl[\frac{b}{(x_{1}\cdots x_{d})^{p^{n}}f}\Bigr].$
Then for some $c\in\overline{A}$,
$(a,0,\ldots,0)\equiv(0,\ldots,0,f^{s_{n}}b+f^{s_{n}+1}c)\pmod{(x_{1},\ldots,x_{d})\,Q_{f,\left(s_{0},\ldots,s_{n-1},p\right)}}$.
Hence there exists
$\alpha\in\operatorname{Ker}\!\Bigl(\overline{A}\oplus F_{*}\overline{A}\oplus\cdots\oplus F^{n}_{*}\overline{A}\to Q_{A,f,\left(0,s_{1},\ldots,s_{n}+1\right)}\Bigr)$
such that
$(a,0,\ldots,0)+\alpha\equiv(0,\ldots,0,f^{s_{n}}b+f^{s_{n}+1}c)\pmod{(x_{1},\ldots,x_{d})(\overline{A}\oplus\cdots\oplus F^{n}_{*}\overline{A})}$.
Let $(B,(d))$ be a perfect prism, $\varphi\colon A\to B$ a morphism of $\delta$-rings, and $y\in B$ as in Lemma 4.4. By Lemma 4.3 (using $s_{n}+1\leq p$) we have
$(a,0,\ldots,0)\,y^{1-(s_{n}+1)/p^{n}}$ $\equiv(0,\ldots,0,f^{s_{n}}b+f^{s_{n}+1}c)\,y^{1-(s_{n}+1)/p^{n}}$
$\qquad\pmod{(x_{1},\ldots,x_{d},f)\,Q_{B,d,\left(s_{0},\ldots,s_{n-1},1\right)}}$.
Therefore,
$V(\tau^{\prime}_{n})$ $=V\!\Bigl(\Bigl[\frac{b}{(x_{1}\cdots x_{d})^{p^{n}}f}\Bigr]\Bigr)=V\!\Bigl(\Bigl[\frac{f^{s_{n}}b}{(x_{1}\cdots x_{d}f)^{p^{n}}}\Bigr]y^{p^{n}-s_{n}-1}\Bigr)$
$=\Bigl[\frac{(0,\ldots,0,f^{s_{n}}b)}{x_{1}\cdots x_{d}f}\Bigr]\,y^{1-(s_{n}+1)/p}=\Bigl[\frac{(a,0,\ldots,0)}{x_{1}\cdots x_{d}f}\Bigr]\,y^{1-(s_{n}+1)/p}$
$=\Phi_{B,d,\left(s_{0},\ldots,s_{n-1},1\right)}\!\Bigl(\Bigl[\frac{a}{x_{1}\cdots x_{d}f}\Bigr]y^{1-(s_{n}+1)/p^{n}}\Bigr)=\Phi_{B,d,\left(s_{0},\ldots,s_{n-1},1\right)}(\tau^{\prime}y^{1-(s_{n}+1)/p^{n}}),$
where $\tau^{\prime}=(H^{d+1}_{\mathfrak{m}}(\overline{A})\to H^{d+1}_{\mathfrak{m}}(\overline{B/(d)}))(\tau)$ and $\tau^{\prime}_{n}$ is defined similarly.
Since $B/(d)$ is Cohen–Macaulay, the map
$V\colon H^{d+1}_{\mathfrak{m}}(\overline{B/(d)})\to H^{d+1}_{\mathfrak{m}}(Q_{B,d,\left(s_{0},\ldots,s_{n-1},1\right)})$
is injective; and because $A\to B/(d)$ is pure, $\tau^{\prime}y^{1-(s_{n}+1)/p^{n}}\neq 0$. As $s_{n}+1\leq p$, this implies $\tau^{\prime}y^{1-1/p^{n-1}}\neq 0$ for all $n$. Since $y\equiv f\pmod{(d)}$, the natural map
$H^{d}_{\mathfrak{m}}(\overline{A/f})\longrightarrow H^{d}_{\mathfrak{m}}(\overline{B/(d,y^{1/p^{\infty}})})$
is injective; hence $A/f$ is perfectoid pure.
Next, we prove (2). For $n\geq 1$ set
$\alpha_{n}$ $=\sum_{i=1}^{n}\frac{p-1-s_{i}}{p^{i}},$ $\qquad\alpha$ $=\lim_{n\to\infty}\alpha_{n},$
$R$ $=A/f,$ $\qquad S$ $=B/(y^{1/p^{\infty}}),$ $\qquad R_{\infty}$ $=B/(d,y^{1/p^{\infty}}).$
Let $\eta^{\prime}=(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R_{\infty}}))(\eta)$. By Proposition 2.2,
$\operatorname{ppt}(R,p)=\sup\{\beta\in\mathbb{Z}[1/p]_{\geq 0}\mid p^{\beta}\eta^{\prime}\neq 0\}.$
As in (1), we have
$\Phi_{B,d,\left(s_{0},\ldots,s_{n-1},1\right)}\bigl(\tau^{\prime}y^{1-1/p^{\,n-1}}\bigr)\neq 0.$
Applying Proposition 4.2, we obtain
$p^{\alpha_{n-1}+\frac{p-1}{p^{n}}}\ \tau^{\prime}y^{1-1/p^{\,n-1}}\neq 0\quad\text{in }H^{d+1}_{\mathfrak{m}}(\overline{B/(d)})\ \text{for all }n\geq 1.$
Fix $n\geq 2$. Then for every $n^{\prime}\geq n$, $p^{\alpha_{n-1}}\tau^{\prime}y^{1-1/p^{\,n^{\prime}-1}}\neq 0$, hence $p^{\alpha_{n-1}}\eta^{\prime}\neq 0$. Thus $\operatorname{ppt}(R,p)\geq\alpha_{n-1}$ for all $n\geq 2$, and so $\operatorname{ppt}(R,p)\geq\alpha$.
For the reverse inequality, functoriality (Proposition 3.4) gives the commutative square
(4.1) ${\overline{R}}$${Q_{R,f,\left(s_{0},\ldots,s_{n}\right)}}$${\overline{R_{\infty}}}$${Q_{S,d,\left(s_{0},\ldots,s_{n}\right)}\,.}$$\scriptstyle{\Phi_{A,f,\left(s_{0},\ldots,s_{n}\right)}}$$\scriptstyle{\Phi_{S,d,\left(s_{0},\ldots,s_{n}\right)}}$
By the definition of $\boldsymbol{s}(f)$, $\Phi_{A,f,\left(s_{0},\ldots,s_{n}\right)}(\eta)=0$; hence $\Phi_{S,d,\left(s_{0},\ldots,s_{n}\right)}(\eta^{\prime})=0$. By Proposition 4.2, $p^{\alpha_{n}+\frac{1}{p^{n}}}\eta^{\prime}=0$. Therefore $\operatorname{ppt}(R,p)\leq\alpha_{n}+\tfrac{1}{p^{n}}$ for all $n\geq 1$, and passing to the limit yields $\operatorname{ppt}(R,p)\leq\alpha$. This proves (2). ∎
Question 4.6.
Does the converse direction of Theorem 4.5 hold? That is, let $(A,\mathfrak{m})$ be a regular local ring admitting a finite Frobenius lift $\phi$, and suppose $p\in\mathfrak{m}$. Fix the $\delta$-ring structure on $A$ induced by $\phi$. Let $f\in A$ be such that $p,f$ form a regular sequence. If $A/f$ is perfectoid pure, do we have $s_{i}\leq p-1$ for all $i\geq 1$?
Theorem 4.7.
(Theorem B) Let $(A,\mathfrak{m})$ be a regular local ring with a Frobenius lift $\phi$ and $p\in\mathfrak{m}$. Fix the $\delta$-ring structure on $A$ induced by $\phi$. Let $f\in A$ be such that $p,f$ form a regular sequence. Let $\boldsymbol{s}(f)=(s_{i})_{i\geq 0}$ be the splitting–order sequence of $f$.
  1. (1)
    If $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=\frac{p-2}{p-1}$, then $s_{n}=1$ for every $n\geq 1$.
  2. (2)
    If $s_{1}=\cdots=s_{r}=p-1$ and $s_{r+1}=p$ for some $r\geq 2$, then $A/f$ is not perfectoid pure.
  3. (3)
    If $A/f$ is regular, then $s_{n}\leq p-1$ for every $n\geq 0$.
Proof.
First, we prove (1). We use the same notation as in the proof of Theorem 4.5, and prove the claim by induction on $n$.
We note that
$\frac{p-2}{p-1}=\sum_{i\geq 1}\frac{p-2}{p^{i}}.$
Assume that $s_{i}=1$ for $n-1\geq i\geq 1$. Then by (4.1) and the argument in the proof of Theorem 4.5, we have
$p^{\frac{p-2}{p}+\cdots+\frac{p-2}{p^{n-1}}+\frac{p-s_{n}}{p^{n}}}\,\eta^{\prime}=0.$
Hence
$\operatorname{ppt}(A/f,p)=\frac{p-2}{p-1}\leq\frac{p-2}{p}+\cdots+\frac{p-2}{p^{n-1}}+\frac{p-s_{n}}{p^{n}}.$
This inequality implies $s_{n}\leq 1$.
If $s_{n}=0$, then $A/f$ is quasi-$F$-split of height $n$ by Proposition 3.8, and thus $\operatorname{ppt}(A/f,p)>\frac{p-2}{p-1}$ by [Yoshikawa25]*Theorem A, a contradiction. Therefore $s_{n}=1$, as desired.
Next, we prove (2). By a diagram (4.1) and $\Phi_{R,f,\left(s_{0},s_{1},\ldots,s_{r+1}\right)}(\eta)=0$, we obtain
$\Phi_{S,d,\left(s_{0},s_{1},\ldots,s_{r+1}\right)}(\eta^{\prime})=0.$
Since $s_{1}=\cdots=s_{r}=p-1$ and $s_{r+1}=p$, we have $\eta^{\prime}=0$ by Proposition 4.2, as desired.
Finally, we prove (3). If $\overline{R}$ is regular, then $s_{n}=0$ for every $n\geq 0$ by Proposition 3.8. Hence we assume that $\overline{R}$ is not regular; in particular, $f\in(\mathfrak{m}^{2},p)$. Write $f\equiv pv\pmod{\mathfrak{m}^{2}}$. Since $f\notin\mathfrak{m}^{2}$, we have $v\notin\mathfrak{m}$. It follows that
$\Delta(f)\equiv v^{p}\pmod{(p,\mathfrak{m})}$,
so $\Delta(f)$ is a unit. Therefore $f$ is a distinguished element in $B$, and in particular we have $fS=dS$.
Since the map $A/f\to S/d$ is pure, it follows that $A/f^{s}\to S/d^{s}$ is pure for every $s\geq 1$. By the exact sequence in Proposition 3.5, the induced map
$Q_{A,f,\left(s_{0},\ldots,s_{n}+1\right)}\longrightarrow Q_{S,d,\left(s_{0},\ldots,s_{n}+1\right)}$
is pure for all $n\geq 1$.
For $n\geq 1$, set
$\alpha_{n}:=\frac{p-1-s_{1}}{p}+\cdots+\frac{p-1-s_{n}}{p^{n}}.$
We prove by induction on $n\geq 1$ that
$p^{\alpha_{n}}\eta^{\prime}\neq 0,\qquad p^{\alpha_{n}+1/p^{n}}\eta^{\prime}=0,\qquad\text{and}\qquad s_{n}\leq p-1.$
Assume $s_{i}\leq p-1$ for all $i\leq n-1$. By the definition of $s_{n}$, we have
$\Phi_{A,f,\left(s_{0},\ldots,s_{n}\right)}(\eta)=0,$
and hence
$\Phi_{S,d,\left(s_{0},\ldots,s_{n}\right)}(\eta^{\prime})=0.$
By Proposition 4.2, this implies
$p^{\alpha_{n}+1/p^{n}}\eta^{\prime}=0.$
If $n=1$, the perfectoid purity of $R$ gives $s_{1}\leq p-1$. If $n\geq 2$ and $s_{n}=p$, then
$p^{\alpha_{n-1}}\eta^{\prime}=0,$
contradicting the induction hypothesis. Therefore $s_{n}\leq p-1$, and moreover
$\Phi_{A,f,\left(s_{0},\ldots,s_{n}+1\right)}(\eta)\neq 0.$
Since the map
$Q_{A,f,\left(s_{0},\ldots,s_{n}+1\right)}\longrightarrow Q_{S,d,\left(s_{0},\ldots,s_{n}+1\right)}$
is pure, we obtain
$\Phi_{S,d,\left(s_{0},\ldots,s_{n}+1\right)}(\eta^{\prime})\neq 0.$
By Proposition 4.2, it follows that
$p^{\alpha_{n}}\eta^{\prime}\neq 0,$
as desired. ∎
Proof of Theorem B.
Assertions (2) and (3) follow from Theorem 4.7. Thus, we may assume that $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)\geq\frac{p-2}{p-1}$. If equality holds, the claim follows from Theorem 4.7(1). If instead $\operatorname{ppt}(A/f,p)>\frac{p-2}{p-1}$, then $A/f$ is quasi-$F$-split by [Yoshikawa25]*Theorem A. By Proposition 3.8, we obtain $s_{n}\leq 1$ for every $n\geq 1$, as desired. ∎
5. A computational method for splitting-order sequences
In this section, we provide a computational criterion for determining the splitting-order sequence, analogous to Fedder’s criterion in $F$-singularity theory. Using this, we give an explicit connection between the perfectoid $p$-purity threshold and the $F$-pure threshold in the regular case.
Notation 5.1.
Let $(A,\mathfrak{m})$ be a regular local ring equipped with a Frobenius lift $\phi$, and assume that $p\in\mathfrak{m}$. Fix the $\delta$-ring structure on $A$ induced by $\phi$. Let $f\in A$ be such that $p,f$ form a regular sequence, and let $\boldsymbol{s}(f)=(s_{i})_{i\geq 0}$ denote the splitting–order sequence of $f$. We set $\mathfrak{m}^{[p^{e}]}:=(p,x^{p^{e}}\mid x\in\mathfrak{m})$.
Fix a generator $u\in\operatorname{Hom}_{\overline{A}}(F_{*}\overline{A},\overline{A})$. Note that since $\phi$ is flat, the Frobenius $F$ on $\overline{A}$ is also flat.
For $n\geq 1$, set
$Q_{n}:=F_{*}\overline{A}\oplus F^{2}_{*}\overline{A}\oplus\cdots\oplus F^{n}_{*}\overline{A}.$
Then we have a natural isomorphism
(5.1) $\operatorname{Hom}_{A}(Q_{n},\overline{A})\;\simeq\;\operatorname{Hom}_{A}(F^{n}_{*}\overline{A},\overline{A})\oplus\cdots\oplus\operatorname{Hom}_{A}(F_{*}\overline{A},\overline{A})\;\overset{(\star_{1})}{\simeq}\;F_{*}^{n}\overline{A}\oplus\cdots\oplus F_{*}\overline{A},$
where $(\star_{1})$ is given, for each $1\leq e\leq n$, by
$F^{e}_{*}\overline{A}\longrightarrow\operatorname{Hom}_{A}(F^{e}_{*}\overline{A},\overline{A}),\qquad F^{e}_{*}a\longmapsto\bigl(F^{e}_{*}b\longmapsto u^{e}(F^{e}_{*}(ab))\bigr).$
The homomorphism corresponding to $(F^{n}_{*}g_{n},\ldots,F_{*}g_{1})$ under (5.1) is denoted by $\psi_{(g_{1},\ldots,g_{n})}$.
Finally, for integers $0\leq l_{1},\ldots,l_{n-1}\leq p-1$ and $0\leq l_{n}\leq p$, we define ideals $I(l_{1},\ldots,l_{n})$ of $\overline{A}$ inductively by
  • $I(l_{n}):=f^{p-l_{n}}\overline{A}$,
  • once $I(l_{2},\ldots,l_{n})$ is defined, set
    $I(l_{1},\ldots,l_{n}):=f^{p-l_{1}-1}\,u\bigl(F_{*}(\Delta(f)^{l_{1}}I(l_{2},\ldots,l_{n}))\bigr)+f^{p-l_{1}}\overline{A}.$
The proof of Theorem 5.2 is almost identical to a proof of [Yoshikawa25]*Theorem 4.13, however, for the convenience of the reader, we provide a proof.
Theorem 5.2.
With notation as in 5.1. For $n\geq 1$, if $s_{1},\ldots,s_{n-1}\leq p-1$, then
$s_{n}=\max\bigl\{\,0\leq s\leq p\,\bigm|\,I(s_{1},\ldots,s_{n-1},s)\subseteq\mathfrak{m}^{[p]}\,\bigr\}.$
Proof.
Fix $n\geq 1$, $0\leq l_{1},\ldots,l_{n-1}\leq p-1$, and $0\leq l_{n}\leq p$. Consider the composition
$\Phi\colon\overline{R}\xrightarrow{\Phi_{f,\left(0,l_{1},\ldots,l_{n}\right)}}Q_{f,\left(0,l_{1},\ldots,l_{n}\right)}\xrightarrow{\pi}F_{*}Q_{f,\left(l_{1},\ldots,l_{n}\right)},$
where $\pi(a_{0},\ldots,a_{n})=(a_{0}^{p}+a_{1},\ldots,a_{n})$; note that $\pi$ is an isomorphism.
The map $\Phi$ induces
$\operatorname{Hom}_{\overline{R}}(Q_{f,\left(l_{1},\ldots,l_{n}\right)},\overline{R})\longrightarrow\overline{R},\qquad\psi\longmapsto\psi\circ\Phi(1)=\psi(1,0,\ldots,0).$
It suffices to show that
(5.2) $u(F_{*}I(l_{1},\ldots,l_{n}))\,\overline{R}=\mathrm{Im}\!\bigl(\operatorname{Hom}_{\overline{R}}(Q_{f,\left(l_{1},\ldots,l_{n}\right)},\overline{R})\to\overline{R}\bigr).$
Step 1. Proof of the inclusion $(\supseteq)$ in (5.2). Take $\psi^{\prime}\in\operatorname{Hom}_{\overline{R}}(Q_{f,\left(l_{1},\ldots,l_{n}\right)},\overline{R})$. Since there is a surjection $Q_{n}\twoheadrightarrow Q_{f,\left(l_{1},\ldots,l_{n}\right)}$ and $Q_{n}$ is a free $\overline{A}$-module, there exists an $\overline{A}$-linear map $\psi\colon Q_{n}\to\overline{A}$ such that
(5.3) $\psi(a_{1}f^{l_{1}},-a_{1}^{p}\Delta(f)^{l_{1}}+a_{2}f^{l_{2}},\ldots,-a_{n-1}^{p}\Delta(f)^{l_{n-1}}+a_{n}f^{l_{n}})\in f\overline{A}$
and such that $(\overline{A}\to\overline{R})\bigl(\psi(1,0,\ldots,0)\bigr)=\psi^{\prime}(1,0,\ldots,0)$.
Hence it suffices to show that $\psi(1,0,\ldots,0)\in u(F_{*}I(l_{1},\ldots,l_{n}))+f\overline{A}$.
Write $\psi=\psi_{(g_{1},\ldots,g_{n})}$ for some $g_{1},\ldots,g_{n}\in\overline{A}$. From (5.3), taking $a_{0}=\cdots=a_{n-1}=0$, we have $u^{n}(F^{n}_{*}(g_{n}af^{l_{n}}))\in f\overline{A}$ for all $a\in\overline{A}$, hence $g_{n}\in f^{p^{n}-l_{n}}\overline{A}$. Write $g_{n}=h_{n}f^{p^{n}-p}$; then $h_{n}\in I(l_{n})$.
Similarly, for $1\leq i\leq n-1$, we have
$g_{i}f^{l_{i}}-u(F_{*}(g_{i+1}\Delta(f)^{l_{i}}))\in f^{p^{i}}\overline{A}.$
This implies $g_{i}=f^{p^{i}-p}h_{i}$ for some $h_{i}\in I(l_{i},\ldots,l_{n})$, and in particular $g_{1}=h_{1}\in I(l_{1},\ldots,l_{n})$. Therefore,
$\psi(1,0,\ldots,0)=u(F_{*}h_{1})\in u(F_{*}I(l_{1},\ldots,l_{n})),$
as desired.
Step 2. Proof of the inclusion $(\subseteq)$ in (5.2). Take $h_{1}\in I(l_{1},\ldots,l_{n})$. Then there exist elements $h_{2},\ldots,h_{n}$ with $h_{i+1}\in I(l_{i+1},\ldots,l_{n})$ such that
$h_{i}-f^{p-l_{i}-1}u(F_{*}(h_{i+1}\Delta(f)^{l_{i}}))\in f^{p-l_{i}}\overline{A}\qquad(1\leq i\leq n-1).$
Set $g_{i}:=f^{p^{i}-p}h_{i}$. Then for every $a\in\overline{A}$,
$u^{i}(F^{i}_{*}(ag_{i}f^{l_{i}}))-u^{i+1}(F^{i+1}_{*}(a^{p}g_{i+1}\Delta(f)^{l_{i}}))$
$\qquad=u^{i}\!\bigl(F^{i}_{*}(a(g_{i}f^{l_{i}}-u(F_{*}g_{i+1}\Delta(f)^{l_{i}})))\bigr)$
$\qquad=u^{i}\!\bigl(F^{i}_{*}(af^{p^{i}-p+l_{i}}(h_{i}-f^{p-l_{i}-1}u(F_{*}(h_{i+1}\Delta(f)^{l_{i}}))))\bigr)\in f\overline{A}.$
Hence $\psi=\psi_{(g_{1},\ldots,g_{n})}$ induces a homomorphism $\psi^{\prime}\colon Q_{f,\left(l_{1},\ldots,l_{n}\right)}\to\overline{R}$ such that
$\psi^{\prime}(1,0,\ldots,0)=(\overline{A}\to\overline{R})(u(F_{*}h_{1}))\in u(F_{*}I(l_{1},\ldots,l_{n}))\,\overline{R},$
which proves $(\subseteq)$ in (5.2). ∎
Theorem 5.3.
Let $(A,\mathfrak{m})$ be a regular local ring equipped with a Frobenius lift $\phi$, and assume that $p\in\mathfrak{m}$. Fix the $\delta$-ring structure on $A$ induced by $\phi$. Let $f\in A$ be such that $p,f$ form a regular sequence. If $A/f$ is regular, then
$\operatorname{ppt}(A/f,p)=\operatorname{fpt}(\overline{A},\overline{f}),$
where $\overline{f}$ denotes the image of $f$ under $A\to\overline{A}$.
Proof.
Define
$\nu_{f}(p^{e}):=\max\{\,N\mid\overline{f}^{N}\notin\mathfrak{m}^{[p^{e}]}\,\},$
so that $\operatorname{fpt}(\overline{A},\overline{f})=\lim_{e\to\infty}\nu_{f}(p^{e})/p^{e}.$
Let $\boldsymbol{s}(f)=(s_{i})_{i\geq 0}$ be the splitting–order sequence of $f$. By Theorem 4.7(3), we have $s_{i}\leq p-1$ for all $i\geq 0$.
Step 1. Description of $I(l_{1},\ldots,l_{n})$. We claim that for $0\leq l_{1},\ldots,l_{n}\leq p-1$,
(5.4) $I(l_{1},\ldots,l_{n})=u^{n-1}\!\bigl(F^{n-1}_{*}(f^{p^{n}-p^{n-1}l_{1}-p^{n-2}l_{2}-\cdots-l_{n}}\overline{A})\bigr).$
We prove this by induction on $n$.
If $n=1$, the formula is clear. Assume $n\geq 2$ and the statement holds for n1n-1. Since $\Delta(f)$ is a unit by the proof of Theorem 4.7(3), we compute:
$I(l_{1},\ldots,l_{n})$ $=f^{p-l_{1}-1}\,u(F_{*}I(l_{2},\ldots,l_{n}))+f^{p-l_{1}}\overline{A}$
$\overset{(\star_{1})}{=}u^{n-1}\!\bigl(F^{n-1}_{*}(f^{p^{n}-p^{n-1}l_{1}-\cdots-l_{n}}\overline{A})\bigr)+f^{p-l_{1}}\overline{A}$
$\overset{(\star_{2})}{=}u^{n-1}\!\bigl(F^{n-1}_{*}(f^{p^{n}-p^{n-1}l_{1}-\cdots-l_{n}}\overline{A})\bigr),$
where $(\star_{1})$ follows from the induction hypothesis, and $(\star_{2})$ follows from
$u^{n-1}\!\bigl(F^{n-1}_{*}(f^{p^{n}-p^{n-1}l_{1}-\cdots-l_{n}}\overline{A})\bigr)\supseteq u^{n-1}\!\bigl(F^{n-1}_{*}(f^{p^{n}-p^{n-1}l_{1}}\overline{A})\bigr)\supseteq f^{p-l_{1}}\overline{A}.$
This proves (5.4).
Step 2. Relation between $\nu_{f}(p^{n})$ and $\boldsymbol{s}(f)$. We now show that
$\frac{\nu_{f}(p^{n})}{p^{n}}=\sum_{1\leq i\leq n}\frac{p-1-s_{i}}{p^{i}}=:\alpha_{n}.$
By (5.4) and Theorem 5.2, we have
$\overline{f}^{\,p^{n}-p^{n-1}s_{1}-\cdots-s_{n}}\in\mathfrak{m}^{[p^{n}]},\qquad\overline{f}^{\,p^{n}-p^{n-1}s_{1}-\cdots-s_{n}-1}\notin\mathfrak{m}^{[p^{n}]}.$
Hence
$\frac{\nu_{f}(p^{n})}{p^{n}}=\frac{p^{n}-p^{n-1}s_{1}-\cdots-s_{n}-1}{p^{n}}=\alpha_{n}.$
Taking the limit as $n\to\infty$, we obtain
$\operatorname{fpt}(\overline{A},\overline{f})=\lim_{n\to\infty}\alpha_{n}=\operatorname{ppt}(A/f,p),$
where the last equality follows from Theorem 4.5. ∎
Corollary 5.4.
(Theorem C) Let $(R,\mathfrak{m})$ be a regular local ring with $0\neq p\in\mathfrak{m}$ such that $\overline{R}$ is $F$-finite. Then $\operatorname{ppt}(R,p)$ is a rational number.
Proof.
We may assume that $R$ is complete. Let $k$ be the residue field, and let $C$ be a complete discrete valuation field with residue field $k$. Then there exists a ring $A$ of formal power series over $C$ and an element $f\in A$ such that $A/f\simeq R$. Since the residue field of $C$ is $F$-finite, AA has a finite Frobenius lift. By Theorem 5.3, we obtain
$\operatorname{ppt}(R,p)=\operatorname{fpt}(\overline{A},\overline{f}),$
and the right-hand side is rational by [BMS]*Theorem 3.1. ∎
6. Examples
Finally, we present explicit examples illustrating how the splitting-order sequence and perfectoid purity behave in concrete settings. These computations demonstrate the rationality and stability phenomena established in the previous sections.
Notation 6.1.
With the notation as in 5.1, we assume
$A:=W(k)[[x_{1},\ldots,x_{N}]]$
for a perfect field $k$ and $\phi(x_{i})=x_{i}^{p}$ for $1\leq i\leq N$. We set $\overline{x}_{i}:=(A\to\overline{A})(x_{i})$ for $1\leq i\leq N$. Since
$\{F_{*}(\overline{x}_{1}^{i_{1}}\cdots\overline{x}_{N}^{i_{N}})\mid 0\leq i_{1},\ldots,i_{N}\leq p-1\}$
is a basis of $F_{*}\overline{A}$ over $\overline{A}$, we can take a generator $u\in\operatorname{Hom}_{\overline{A}}(F_{*}\overline{A},\overline{A})$ as an element of dual basis corresponding to $F_{*}(\overline{x}_{1}^{p-1}\cdots\overline{x}_{N}^{p-1})$.
Theorem 6.2.
With the notation as in 6.1, let $f=x_{1}^{N}+\cdots+x_{N}^{N}$ and assume $p>N$. For each $e\geq 1$, take an integer $0\leq s_{e}\leq N-2$ satisfying $s_{e}+1\equiv p^{e}\pmod{N}$. Then the splitting–order sequence of $f$ is $\boldsymbol{s}(f)=(0,s_{1},s_{2},\ldots)$. In particular, $A/f$ is perfectoid pure.
Proof.
We set $X:=x_{1}\cdots x_{N}$ and prove that
$I(s_{1},\ldots,s_{n})\subseteq\mathfrak{m}^{[p]}\quad\text{and}\quad I(s_{1},\ldots,s_{n}+1)\nsubseteq\mathfrak{m}^{[p]}$
for all $n\geq 1$ by induction on $n$.
For $n=1$, since
$f^{p-s_{1}-1}\equiv X^{p-s_{1}-1}\pmod{\mathfrak{m}^{[p]}}$,
we obtain $f^{p-s_{1}}\in\mathfrak{m}^{[p]}$, as desired.
Next, we fix $n\geq 2$ and assume the claim holds for all smaller $n$. We show that
(6.1) $I(s_{1},\ldots,s_{n-1},s)\equiv u^{e}\bigl(F^{e}_{*}(X^{p^{e+1}-p(s_{e}+1)}\Delta(f)^{s_{e}}I(s_{e+1},\ldots,s_{n-1},s))\bigr)\pmod{\mathfrak{m}^{[p]}}$
for every $1\leq e\leq n-1$ and $0\leq s\leq p$, by induction on $e$.
For $e=1$, the equality follows directly from the definition of $I(\cdot)$ and the congruence $f^{p-s_{1}-1}\equiv X^{p-s_{1}-1}\pmod{\mathfrak{m}^{[p]}}$. Suppose $e\geq 2$, and assume that (6.1) holds for e1e-1. We claim that
(6.2) $\Delta(f)^{s_{e-1}}f^{p-s_{e}-1}\equiv X^{p(s_{e-1}+1)-(s_{e}+1)}\pmod{(x_{1}^{pN},\ldots,x_{N}^{pN})}$.
Indeed, since $f=x_{1}^{N}+\cdots+x_{N}^{N}$, we have
$\Delta(f)^{s_{e-1}}f^{p-s_{e}-1}$ $=\frac{1}{p^{s_{e-1}}}(f^{p}-(x_{1}^{pN}+\cdots+x_{N}^{pN}))^{s_{e-1}}f^{p-s_{e}-1}$
$\equiv\frac{1}{p^{s_{e-1}}}f^{p(s_{e-1}+1)-(s_{e}+1)}\pmod{(x_{1}^{pN},\ldots,x_{N}^{pN})}$.
By the definition of $s_{e}$, we have $p(s_{e-1}+1)\equiv s_{e}+1\pmod{N}$, and the coefficient of $X^{p(s_{e-1}+1)-(s_{e}+1)}$ in $f^{p(s_{e-1}+1)-(s_{e}+1)}$ is nonzero and not divisible by $p^{s_{e-1}+1}$. Thus, (6.2) holds.
It follows that
(6.3) $X^{p^{e}-p(s_{e-1}+1)}\Delta(f)^{s_{e-1}}f^{p-s_{e}-1}\equiv X^{p^{e}-(s_{e}+1)}\pmod{\mathfrak{m}^{[p^{e}]}}$
and $X^{p^{e}-p(s_{e-1}+1)}\Delta(f)^{s_{e-1}}f^{p-s_{e}}\in\mathfrak{m}^{[p^{e}]}$. Substituting (6.3) into the definition of $I(\cdot)$, we obtain
$I(s_{1},\ldots,s_{n})$ $\equiv u^{e-1}\bigl(F^{e-1}_{*}(X^{p^{e}-p(s_{e-1}+1)}\Delta(f)^{s_{e-1}}I(s_{e},\ldots,s_{n}))\bigr)$
$\equiv u^{e-1}\bigl(F^{e-1}_{*}(X^{p^{e}-(s_{e}+1)}u(F_{*}(\Delta(f)^{s_{e}}I(s_{e+1},\ldots,s_{n}))))\bigr)$
$=u^{e}\bigl(F^{e}_{*}(X^{p^{e+1}-p(s_{e}+1)}\Delta(f)^{s_{e}}I(s_{e+1},\ldots,s_{n}))\bigr)\pmod{\mathfrak{m}^{[p]}}$,
which proves (6.1) for all $e$.
Taking e=n1e=n-1 in (6.1), we have
$I(s_{1},\ldots,s_{n-1},s)$ $\equiv u^{n-1}\bigl(F^{n-1}_{*}(X^{p^{n}-p(s_{n-1}+1)}\Delta(f)^{s_{n-1}}f^{p-s}\overline{A})\bigr)$
$\equiv\begin{cases}0&\text{if }s=s_{n},\\[3.0pt] X^{p-1}&\text{if }s=s_{n}+1,\end{cases}\pmod{\mathfrak{m}^{[p]}}$,
and the result follows. ∎
Proposition 6.3.
With the notation as in 6.1, assume that $f\in\mathfrak{m}^{[p]}$.
  1. (1)
    If
    $f^{p-1}\Delta(f)^{p-1}\equiv x_{1}^{p-1}\cdots x_{N}^{p-1}\pmod{\mathfrak{m}^{[p^{2}]}}$,
    then we have $\boldsymbol{s}(f)=(0,p-1,0,p-1,0,\ldots)$. Hence $A/f$ is perfectoid pure, and $\operatorname{ppt}(A/f,p)=\frac{1}{p+1}$ by Theorem 4.5.
  2. (2)
    If $\Delta(f)^{p-1}\in\mathfrak{m}^{[p^{2}]}$, then $A/f$ is not perfectoid pure.
  3. (3)
    If there exists $f^{\prime}\in A$ such that $f=f^{\prime}+px_{1}\cdots x_{N}$ and $\Delta(f^{\prime})\in\mathfrak{m}^{[p^{2}]}$, then $\boldsymbol{s}(f)=(0,p-1,p-1,\ldots)$. Hence $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$ by Theorem 4.5.
Proof.
Proof of (1). We claim that
$J_{2m}:=I(\underbrace{p-1,0,p-1,\ldots,p-1,1}_{2m\text{ terms}})\equiv(x_{1}\cdots x_{N})^{p-1}\overline{A}\pmod{\mathfrak{m}^{[p]}}$
for all $m\geq 1$, and proceed by induction on $m$.
For $m=1$,
$I(p-1,1)$ $=u\bigl(F_{*}(\Delta(f)^{p-1}f^{p-1}\overline{A})\bigr)+f\overline{A}$
$\equiv(x_{1}\cdots x_{N})^{p-1}\overline{A}\pmod{\mathfrak{m}^{[p]}}$.
For $m\geq 2$,
$J_{2m}$ $=u\bigl(F_{*}(\Delta(f)^{p-1}f^{p-1}u(F_{*}J_{2m-2}))\bigr)+f\overline{A}$
$\equiv u\bigl(F_{*}((x_{1}\cdots x_{N})^{p^{2}-1}u(F_{*}J_{2m-2}))\bigr)\overset{(\star_{1})}{\equiv}(x_{1}\cdots x_{N})^{p-1}\pmod{\mathfrak{m}^{[p]}}$,
where $(\star_{1})$ follows from the induction hypothesis.
Next, we show that
$J_{2m-1}:=I(\underbrace{p-1,0,p-1,\ldots,0,p-1}_{(2m-1)\text{ terms}})\subseteq\mathfrak{m}^{[p]}$
for all $m\geq 1$.
For $m=1$,
$J_{1}=I(p-1)=f\overline{A}\subseteq\mathfrak{m}^{[p]}.$
For $m\geq 2$,
$J_{2m-1}$ $=u\bigl(F_{*}(\Delta(f)^{p-1}f^{p-1}u(F_{*}J_{2m-3}))\bigr)+f\overline{A}$
$\equiv u\bigl(F_{*}((x_{1}\cdots x_{N})^{p^{2}-1}u(F_{*}J_{2m-3}))\bigr)\overset{(\star_{2})}{\equiv}0\pmod{\mathfrak{m}^{[p]}}$,
where $(\star_{2})$ follows from the induction hypothesis.
Finally, we prove that
$J^{\prime}_{2m+1}:=I(\underbrace{p-1,0,p-1,\ldots,0,p}_{(2m+1)\text{ terms}})\equiv(x_{1}\cdots x_{N})^{p-1}\pmod{\mathfrak{m}^{[p]}}$
for all $m\geq 1$.
For $m=1$,
$J^{\prime}_{3}=I(p-1,0,p)=u\bigl(F_{*}(\Delta(f)^{p-1}f^{p-1})\bigr)+f\overline{A}\equiv(x_{1}\cdots x_{N})^{p-1}\pmod{\mathfrak{m}^{[p]}}$.
For $m\geq 2$,
$J^{\prime}_{2m+1}$ $=u\bigl(F_{*}(\Delta(f)^{p-1}f^{p-1}u(F_{*}J^{\prime}_{2m-1}))\bigr)+f\overline{A}$
$\equiv u\bigl(F_{*}((x_{1}\cdots x_{N})^{p^{2}-1}u(F_{*}J^{\prime}_{2m-1}))\bigr)\overset{(\star_{3})}{\equiv}(x_{1}\cdots x_{N})^{p-1}\pmod{\mathfrak{m}^{[p]}}$,
where $(\star_{3})$ follows from the induction hypothesis.
Combining the above computations, we obtain
$\boldsymbol{s}(f)=(0,p-1,0,p-1,0,\ldots),$
as claimed.
Proof of (2). Since $f\in\mathfrak{m}^{[p]}$, we have $s_{1}=p-1$. Moreover,
$I(p-1,p)=u(F_{*}\Delta(f)^{p-1})+f\overline{A}\subseteq\mathfrak{m}^{[p]}.$
Hence $s_{2}=p$. By Theorem 4.7(2), it follows that $A/f$ is not perfectoid pure.
Proof of (3). Note that
$\Delta(f)=\Delta(f^{\prime})-(x_{1}\cdots x_{N})^{p}\equiv-(x_{1}\cdots x_{N})^{p}\pmod{\mathfrak{m}^{[p^{2}]}}$.
We claim that
$J_{m}:=I(\underbrace{p-1,p-1,\ldots,p-1,p}_{m\text{ terms}})\equiv(x_{1}\cdots x_{N})^{p-1}\overline{A}\pmod{\mathfrak{m}^{[p]}}$
for all $m\geq 2$, by induction on $m$.
For $m=2$,
$I(p-1,p)$ $=u\bigl(F_{*}(\Delta(f)^{p-1}\overline{A})\bigr)+f\overline{A}$
$\equiv u\bigl(F_{*}((x_{1}\cdots x_{N})^{p(p-1)}\overline{A})\bigr)\equiv(x_{1}\cdots x_{N})^{p-1}\overline{A}\pmod{\mathfrak{m}^{[p]}}$.
For $m\geq 3$,
$J_{m}$ $=u\bigl(F_{*}(\Delta(f)^{p-1}J_{m-1})\bigr)+f\overline{A}$
$\equiv u\bigl(F_{*}((x_{1}\cdots x_{N})^{p(p-1)}J_{m-1})\bigr)\overset{(\star_{4})}{\equiv}(x_{1}\cdots x_{N})^{p-1}\pmod{\mathfrak{m}^{[p]}}$,
where $(\star_{4})$ follows from the induction hypothesis.
Next, we show that
$J^{\prime}_{m}:=I(\underbrace{p-1,p-1,\ldots,p-1}_{m\text{ terms}})\subseteq\mathfrak{m}^{[p]}$
for all $m\geq 1$.
For $m=1$,
$J^{\prime}_{1}=I(p-1)=f\overline{A}\subseteq\mathfrak{m}^{[p]}.$
For $m\geq 2$,
$J^{\prime}_{m}$ $=u\bigl(F_{*}(\Delta(f)^{p-1}J^{\prime}_{m-1})\bigr)+f\overline{A}$
$\equiv u\bigl(F_{*}((x_{1}\cdots x_{N})^{p(p-1)}J^{\prime}_{m-1})\bigr)\overset{(\star_{5})}{\equiv}0\pmod{\mathfrak{m}^{[p]}}$,
where $(\star_{5})$ follows from the induction hypothesis.
Combining these, we obtain
$\boldsymbol{s}(f)=(0,p-1,p-1,p-1,p-1,\ldots),$
and therefore $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$, as claimed. ∎
Example 6.4.
Let $N=p$ and $f=x_{1}^{p}+\cdots+x_{p}^{p}$, so that $f\in\mathfrak{m}^{[p]}$. Set $f^{\prime}:=f-px_{1}\cdots x_{p}$. Then we have
$\Delta(f^{\prime})=\Delta(f)+(x_{1}\cdots x_{p})^{p}\in\mathfrak{m}^{[p^{2}]}.$
Hence, by Proposition 6.3(3), the quotient $A/f$ is perfectoid pure, and
$\operatorname{ppt}(A/f,p)=0.$
Example 6.5.
Let $N=p+1$, and $f=x_{1}^{p+1}+\cdots+x_{p+1}^{p+1}$. Then $f\in\mathfrak{m}^{[p]}$ and
$f^{p-1}\Delta(f)^{p-1}\equiv(x_{1}\cdots x_{p+1})^{p^{2}-1}\quad\pmod{\mathfrak{m}^{[p]}}$.
Thus, we have $\boldsymbol{s}(f)=(0,p-1,0,p-1,0,\ldots)$ and $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=\frac{1}{p+1}$ by Proposition 6.3 (1).
Example 6.6.
Let $p=2$.
  1. (1)
    Let $N=5$ and $f=x_{1}^{5}+\cdots+x_{5}^{5}$. Then $A/f$ is not perfectoid pure, while $A/(f+2x_{1}\cdots x_{5})$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$ by Proposition 6.3(2) and (3). Indeed, $f\in\mathfrak{m}^{[p]}$ and $\Delta(f)\in\mathfrak{m}^{[p^{2}]}$.
  2. (2)
    Let $N=4$ and
    $f=x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+x_{4}^{4}+x_{1}^{2}x_{2}^{2}+x_{1}^{2}x_{3}^{2}+x_{2}^{2}x_{3}^{2}+x_{1}x_{2}x_{3}(x_{1}+x_{2}+x_{3}).$
    Then $A/f$ is not perfectoid pure, whereas $A/(f+2x_{1}\cdots x_{4})$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=0$ Proposition 6.3(2) and (3).. Indeed, $f\in\mathfrak{m}^{[p]}$ and $\Delta(f)\in\mathfrak{m}^{[p^{2}]}$.
Example 6.7.
(Fermat K3 surface) Let $N=4$ and $f=x^{4}+y^{4}+z^{4}+w^{4}$, where $x:=x_{1}$, $y:=x_{2}$, $z:=x_{3}$, and $w:=x_{4}$.
  1. (1)
    If $p=2$, then $f\in\mathfrak{m}^{[p]}$ and $\Delta(f)\in\mathfrak{m}^{[p^{2}]}$. Hence $A/f$ is not perfectoid pure, while $A/(f+pxyzw)$ is perfectoid pure by Proposition 6.3 (3).
  2. (2)
    If $p\equiv 1\pmod{4}$, then $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=1$, since $\overline{A/f}$ is $F$-pure.
  3. (3)
    If $p\equiv 3\pmod{4}$, then $A/f$ is perfectoid pure with $\operatorname{ppt}(A/f,p)=\frac{2}{p^{2}-1}$. Indeed, for $p=3$, this follows from Example 6.5. If $p>3$, then by Theorem 6.2 we have $\boldsymbol{s}(f)=(0,2,0,2,\ldots)$, and thus $A/f$ is perfectoid pure with the desired value $\operatorname{ppt}(A/f,p)=\frac{2}{p^{2}-1}$.